### v1

# Large Language User Interfaces: Voice Interactive User Interfaces Powered by LLMs

Abstract
The recent meteoric advancements in large language models have showcased a remarkable capacity for logical reasoning and “comprehension”. These newfound capabilities have opened the door to a new generation of software, as has been made obvious through the innumerable ways they are being applied in the industry. This research focuses on harnessing and guiding the upgraded power of LLMs to construct a framework that can serve as an intermediary between a user and their user interface. By comprehending a user’s needs through a thorough analysis of natural textual inputs, an effectively crafted LLM engine can classify the most likely available application, identify the desired UI component and subsequently execute the user’s expected actions. This integration can evolve static UI systems into highly dynamic and adaptable solutions, introducing a new frontier of intelligent and responsive user experiences. Such a framework can fundamentally shift how users accomplish daily tasks, skyrocket efficiency, and greatly reduce cognitive load.
1 Introduction
The modern world relies on and is driven by software. Embedded systems, command-line interfaces and user interface (UI) software are present across systems all around the world. The ease of use coupled with their intuitive nature has allowed for UI systems to become a staple as a crucial tool in modern software and beyond. UI systems serve as a visually appealing packaging of function calls and event handlers, allowing for complex event pipelines and data flows to be abstracted by buttons, text fields, menus, etc.
The evolutions made in large language models (LLMs) over the past year have exhibited true "cognitive" potential. This potent ability has unveiled innumerable new opportunities to revolutionize the way our contemporary software systems are expected to operate. In this paper, we explore our vision and progress toward developing a UI architectural paradigm which employs a multimodal engine powered by LLMs and state-of-the-art transformer models. This framework aims to abstract monotonous UI interactions with prompting mechanisms that serve as "cognitively aware", powering automated functional calling and data flow pipelines, which translate to full speech-based intelligence control over visual UI systems.
Our approach posits a novel architecture and methodology for UI systems, incorporating logically sound LLM engines capable of interpreting a user’s desired action. This facilitates a synergistic interaction between the software and the user, enabling a more intuitive and user-centred experience, rather than requiring the user to conform to the software’s predefined structures.
2 Motivations of the Research
The catalyst for this research is grounded in the critical examination of current user interface systems, which, while omnipresent and essential, do not come without limitations. Traditional implementations of UIs often require the user to adhere to the methods and system constraints established by the developer and their design. This means that all UI components in an application have specific triggers and their invocation is limited to a strict sequence of user inputs within the applications’ available features. Our paradigm would relieve the user of such restricted, stringent and static rule sets when utilizing software, fundamentally changing the landscape of how applications are developed, used and integrated across all sectors. It would allow for real-time intelligent and efficient control over complex tasks, enabling users to achieve substantially more, in significantly less time.
Our research aims to foster this innovation by bridging the gap that exists between the research and practical integration of LLMs into UI systems. Through development and implementation, valuable empirical insights allow for a demonstration of how LLMs can not only integrate with UI systems but also revolutionize them. From enhancing accessibility for users with disabilities to streamlining complex workflows in enterprise systems, the implications are vast and worthy of exploration. This paper seeks to motivate and propagate a new wave of research and development in this direction, to realize the full potential of LLMs in creating more adaptive, intuitive, and efficient UI systems. This combination of deep theoretical exploration and practical application is vital for the next generation of software development.
3 Challenges
The development of such a framework is not without its challenges. An effective fusion of LLMs and event-driven UIs is a complex task with many critical factors of concern. All classifying, processing, parsing and model inference must be achieved in real-time, the architecture must be capable of scaling and integrating unseen sub-applications in zero-to-few shots, and the framework’s interpretation accuracy of ambiguous user inputs must be top-notch. The greatest challenges encountered during our research and implementation of this framework are as follows:
3.1 Integration Complexity
The underlying challenge across the entire research was effectively fusing LLMs with event-driven UIs. Developing an effective framework capable of being scaled with increasing application counts and fostering potential for implementation as an operating system was always a priority. A paradigm as dynamic as the software it intends to power was vital. Implementing a brute-force hard-coded solution would not serve a great purpose or carry the potential for scalability and impact required for next-generation software. This is why we took our time and carefully crafted each component of this framework with future scaling in mind, leaving room for future advanced integrations such as multi-agent systems as well as more powerful LLM model configurations.
3.2 Annotation Data Structure Selection & Traversal
Another critical point of concern was the establishment of a data structure suitable and most effective for representing the overarching system, applications, as well as all UI components per application. Through our research, a tree structure made for the most accurate representation of the overall system, whilst also allowing for efficient traversal and information retrieval, vital for real-time responses further on.
Figure 1: Visual representation of tree data structure used to store application meta descriptions
The specific implementation of this tree structure still poses room for iterative research concerning scalability, however, regardless of the implementation method, the need for top-down level-by-level comparison is where the tree structure shines. Rather than mapping an input to an application and its UI components through an unstructured or linear search, an organized level-by-level search eliminates redundant checks. Each user input is compared to a tree, which stores textual naming, descriptions, expectations, and examples of each application and its UI components. Our use of a top-down manner comparison between the tree structure and user input ensures that only children of relevant nodes are searched, preserving our engine’s real-time processing goals.
Listing 1: An example of how an application’s component metadata is stored in the program’s node structure
Extract the location: "I am Brian O’Connor residing at Apartment 5A, 654 Peachtree Street in New Town and my email address is brian.oconnor@mailservice.net." || "Address": "Apartment 5A, 654 Peachtree Street in New Town"’,
‘Extract the arithmetic expression: "If you have \$50 and spend \$25, how much money do you have left?" || "\$50 - \$25"’,
Extract the location: "Can you provide me with the current weather conditions in Zurich, Switzerland?" || "Zurich:B-CITY,Switzerland:B-COUNTRY"
3.3 Model Selection & Training
For our framework, an engine that is capable of high-level logical reasoning, understanding and general task capability was paramount. A multimodal pipeline was found to be most efficient in allowing for the balancing of inference times between models with parameter sizes appropriate for the task, as opposed to a single overkill model wasting processing time on smaller classification jobs. Tasks that could be accurately and fastest completed by smaller models were assigned a certain model, while more complex tasks such as input comprehension, parameter extraction and parsing were given their own appropriate models.
For each task, many models were researched, to balance accuracy and speed. We found ROBERTA models (Liu et al. (2019)) fine-tuned on datasets such as SQUAD2 (Rajpurkar et al. (2018)) were efficient at the task. For comprehension and parsing upon our application library, we found that even powerful and large models such as Google’s XL T5 (Raffel et al. (2023)) pre-trained to perform well with general multi-task capabilities simply cannot achieve the desired accuracy across our simple application library. The clear answer would be to employ larger and more powerful models, however, inference times and computational resources quickly become a concern.
Custom fine-tuning was key and opened a pathway for us to research how to achieve generalization well enough to output consistently formatted and precise results. While training models on custom datasets relevant to each application in our library is effective, a scalable and more promising approach would be achieving few/zero-shot capabilities. We are exploring the most effective methods to accomplish this along with scaling our computational resources to open the door to models with larger parameter sizes and more rigorous training.
Table 1: Model Benchmarks
T5
Custom Fine-Tune
T5
Google MTL
BERT
SQuAD2
ELECTRA SQuAD2
Weather 95% 36% 96.8% 76%
Account Form 80% 0% 78% 48%
Simple Calculator 100% 1% 99% 16%
Advanced Calculator 81% 1% 5% 1%
3.4 Training Data Generation & Wrangling
Creating custom datasets with realistic, convoluted and vague prompts reflective of real-life scenarios was tough. We found no available datasets with the level nature of prompts required to be effective for our training. Custom datasets were the best option and were a significant task. This is where generation with top-of-the-line LLMs such as ChatGPT 4 served immense value. Many competing models were tested, however, the adherence to prompt engineering, logical understanding of the task, as well as the depth of versatility GPT 4 (OpenAI (2024)) was able to provide, we found to be unmatched. OpenAI’s API was used for mass data generation and any formatting inconsistencies were fixed by hand. These custom datasets were used to fine-tune LLMs such as varying sizes of Google T5 on various tasks.
4 Proposed LM-UI Integration Framework
With the aforementioned motivations and challenges, we present the LM-UI framework. The integration of a hybrid design that merges an annotated front-end component system with a multimodal classification and logical extraction engine, allows us to achieve a strong logical next-generation user experience (UX). Our framework features a UI paradigm that dispatches its component event handler states to a central Redux store. This store then uses reducers to communicate with the LLM engine, appropriately allowing updates to the corresponding front-end state. This framework allows UI software to interpret and predict a user’s target action, even when provided vague or convoluted prompting.
4.1 Modelling UI Controls
Our approach to fusing LLMs with UI hinges on the modelling of UI components as adaptive entities capable of dynamically responding to natural language inputs from the user. At the core of this integration is a re-imagined method to define UI elements such as buttons, text fields, navigation links, etc. These components are incorporated to be part of an event-driven environment where their behaviours and states are modifiable based on user interactions conveyed through natural language prompts.
4.1.1 Meta Description of UI
Meta descriptions of each UI component present in an application are critical to accuracy in our framework. The use of textual metadata providing detailed specifications of how each UI component fits conceptually within the structure of the application, what its configurable parameters are and examples of potential use cases, prime the application for communication of its expectations and needs with the LLM engine down the road.
4.1.2 Descriptive Component Modeling
Each application is treated as a "node", with all its UI elements being "sub-nodes". These nodes serve as a way to represent the structure of an application. All nodes have textual descriptions outlining their purpose in the application, while sub-nodes have descriptions or example questions related to how they may be prompted for extraction during inference. These details allow our LLM engine to understand and have reference knowledge as to what each application does and wants.
Listing 2: An example of how an application’s component metadata is stored in the program’s node structure
NODES = {
    "Weather": {
      "AppName": "Weather",
      "Description": "A sub-node of Weather. Represents the point where location or query details are input for weather-related inquiries.
    "Parameters": { "City": "What is the location?"}
4.2 Describing Applications as Events & States
To harmonize our front-end applications with our descriptive component modelling, the applications in our framework are modelled upon the concept of events and states. Each UI element has a state for all its configurable parameters. Within each application’s parent component, its states are observed for any changes by React useEffect hooks. When states are updated in response to a user event, the useEffect hook immediately updates the UI to reflect the new states. When the master JSON object that holds the current application indicator along with the states of its respective parameters is updated, it is dispatched to the central store. Each interaction with UI components of an application is an event that is reflected by the redux states and store.
4.2.1 Central Store & Reducers
The Redux store serves as the central point for handling each one of our applications’ states. When a user’s textual prompt is received, it is sent to the LLM engine to be processed. Once the engine returns the parsed result to the root front-end component, it triggers action events. These events dispatch the new states to the Redux central store where reducers are employed. The reducers are functions that take the previous state and an action as arguments and return a new state. They are responsible for handling transitions from one state to another state within the application, ensuring state changes are predictable and transparent. In the context of our framework, when an action is dispatched to the Redux store, the corresponding reducer processes the action, updates the state based on the action type and payload, and returns the updated state. This new state then propagates through the application, triggering UI updates where necessary. The use of reducers thus facilitates a smooth, unidirectional data flow and this separation of concerns ensures that the UI components are kept as stateless as possible, promoting reusability and testability.
4.2.2 Case Study: Application/Task Library
The implementation of this framework serves as a sort of “operating system” that functions over a library of applications. Thus far we have three simple applications in the library: An account sign-up form, a live weather application, and a Calculator. These applications help showcase and validate the core functionality and flexibility of our framework. It is easy to see how this can be scaled up to include various essential applications that would be commonly found on smartphones, tablets, smartwatches, etc. Once we can establish and validate the framework’s potential and ability, more complex and intricate UI applications can be tackled with the appropriate implementation schemes.
4.3 Natural Language Parser
Our LLM engine stands firmly on the foundation established by the preceding components and its ability is solely reliant on the groundwork we laid out. When natural language prompts from the user are sent to the engine, it traverses the following pipeline:
Figure 2: Two-Component Framework
4.3.1 Tokenization
  1. 1.
    An input sentence is split into tokens. BERT uses WordPiece tokenization
  2. 2.
    Special tokens like ‘[CLS]‘ (at the beginning) and ‘[SEP]‘ (at the end) are added
  3. 3.
    The tokenization can be represented as a function that converts a sentence into a sequence of tokens:
[…]
4.3.2 Token to ID Conversion
  1. 1.
    Each token is then mapped to an ID using BERT’s vocabulary. This can be represented as:
    $\text{ID}_{i}=\text{Vocab}(\text{"Token}_{i}\text{")},\text{ for }i=1,2,...,N$
    Where ‘Vocab‘ is a function that returns the ID of a token in the BERT vocabulary.
4.3.3 Segment and Positional Embeddings
  1. 1.
    For BERT, segment embeddings (to differentiate between different sentences) and positional embeddings (to give information about the position of the token in the sequence) are also added
  2. 2.
    The total embedding for a token is then:
    $E_{\text{token}}=E_{\text{word}}(ID)+E_{\text{segment}}+E_{\text{position}}$
  3. 3.
    Here, $E_{\text{word}}(ID)$ is the word embedding corresponding to the token ID, $E_{\text{segment}}$ is the segment embedding, and $E_{\text{position}}$ is the positional embedding.
    Cosine $\text{Cosine Similarity}(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$
The cosine similarity between a user’s input and an annotated node will determine the likelihood that the user desired that node. Whichever node has the highest similarity with the user’s input will end up being the classified application, and the engine will traverse its children nodes as a result. This type of Best-First Search mitigates redundant comparisons through all leaf nodes from incorrect applications, in turn significantly reducing the chance for incorrect matching with incorrect UI components.
Parameter Extraction
Proceeding down the pipeline is the core function of the engine. At this level, a powerful LLM with substantial generalization and multitasking ability is used to compare the annotated children nodes of the classified application and compare each to the user’s input. Through prompt engineering, an LLM such as Google’s T5 model fine-tuned on large text corpora for accurate NLP multitasking abilities, can determine which part of the user’s input corresponds to the current child node being checked. This extraction algorithm functions as a decision-making tool with the use of natural language processing techniques. Its objective is to analyze textual input, categorize it based on predefined criteria, and generate a JSON parameter extraction object, to be relayed to the front end and dispatched to the Redux central store.
Listing 3: Extracted parameters parsed into a JSON object
{"CurrentApp":"AccountForm","Config":{"Name": "Connor Syle", "Address": "34 Coronation Street", "Email": "connor32@outlook.com"}}
4.4 Operational Workflow
  1. 1.
    Input Analysis: The algorithm initially analyzes the input to determine its nature and relevance to each UI component child node under the classified application. This categorization is crucial in deciding the subsequent processing path
  2. 2.
    Conditional Logic: Based on the analysis, the algorithm employs conditional logic to select an appropriate extraction mechanism. Parent nodes of classified applications are checked, to guarantee the most suitable model is employed. This selection process is pivotal to the algorithm’s adaptability and accuracy in parameter extraction.
  3. 3.
    Response Generation: Depending on the categorization:
    • If the input falls into a specific category (e.g., calculations), the algorithm engages a dedicated module designed for that category, employing specialized deep-learning models, such as our custom fine-tuned Google’s T5
    • For other types of input, the algorithm utilizes a different set of models tailored for general language understanding and response formulation
    • This delegation of tasks ensures we use the appropriate model for the task, employing larger more capable models for more demanding tasks such as complex arithmetic extraction, while making use of suitable smaller models for simplistic tasks. This optimization ensures inference time is not wasted on tasks that don’t require it, saving resources along with crucial response time.
  4. 4.
    JSON Output Formatting: Finally, the parameters extracted by the model are formatted into a structured JSON object, optimized and suitable for further processing or display to the end-user once sent back to the client. See Listing 3
This pipeline employed by our LLM engine allows for complex comprehension of textual input from a user, by accurately classifying and extracting parameters that the target application requires. This engine is able to streamline the connection between the user and the software by taking the intentions and requirements of both into account, utilizing its logical reasoning to complete the task as it sees fit. This is a stepping stone towards the right direction in the next generation of software.
5 Evaluation
This section intends to evaluate and validate the abilities of several large language models tasked with serving as the crux of our research. Accuracy in parsed inputs is the driving factor behind this project functioning well. The engine must correctly classify which task the user’s prompt corresponds to, extract, parse, and output the expected parameters in the required JSON format.
5.1 Required Tasks
Classification is accomplished by encoding descriptive nodes for each task along with the input node using an all-MiniLM sentence-transformer model and then selecting the node with the greatest cosine similarity. This classification task is extremely reliable and the chance of failure is observed to be negligible. The task that is critical to evaluate is the parameter extraction once the input’s corresponding task has been established. Extensive experimentation and trials were conducted with plentiful transformer models from the Hugging Face hub. Models specialized in Question Answering as well as Text2Text Generation were found to perform extremely well.
5.2 Model Selection
For this evaluation, we pit the most effective models found during initial testing, with available resources for training, hosting, deployment and inference in mind)—the T5 Text2Text Generator models along with BERT and ELECTRA. The BERT and ELECTRA models were already fine-tuned on the SQuAD 2.0 dataset, which is a popular dataset used in fine-tuning models for logical reasoning and question-answering. On the other hand, the T5 model base and multi-task, trained by Google were also chosen due to their flexibility, scalability, and efficacy in abstraction and multi-task situations.
Figure 3: Accuracy comparison between models engineered for our application’s tasks. Each model was trained on logical and information extraction tasks, and further carefully prompt engineered to achieve the best performance possible for each task. Text2Text Generation Models were fine-tuned on multiple datasets to multitask excellence. The T5 custom model was specifically trained on specialized custom datasets relevant to our application’s tasks.
5.3 Observation
We can see from the following benchmarks in Figure 3 that the custom-trained T5 model does exceptionally well across all tasks, performing the best in three out of four tasks, with the last one being a close tie. The BERT model also performs very well but struggles with the logical capabilities required for the advanced calculator prompts. Google’s multitask-trained model falls extremely short in more logically involved tasks while the ELECTRA model does poorly overall. Further training of the T5 model with more diverse and better quality data would allow it to be the most versatile and powerful model for our project.
GPT Baseline VS Custom Engine Prompt Evaluation
Truth Intention GPT-4 Generated Inputs Multimodal Parser Output Eval
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex P. Globetrotter", "Address": "7890 Nomad Path, Backpack City", "Email": "wanderlust.world@travelbug.com"}}
Oh, before I forget, jot down my email, it’s wanderlust.world@travelbug.com, and people shout out to me as Alex P. Globetrotter when they see me. The address? It’s 7890 Nomad Path, Backpack City. The mailman has it tough, I tell you!
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex P. Globetrotter", "Address": "7890 Nomad Path, Backpack City", "Email": "wanderlust.world@travelbug.com"}}
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex J. Turner", "Address": "768 Rolling Rock Street", "Email": "alex.turns@rocknmail.com"}}
I’m registered under the name Alex J. Turner, but everyone sends their regards to my place at 768 Rolling Rock Street, and for a quicker response, they hit me up at alex.turns@rocknmail.com.
[…]
{"CurrentApp":"AccountForm","Config":{"Name": "Alex J. Turner", "Address": "768 Rolling Rock Street", "Email": "alex.turns@rocknmail.com"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "Amsterdam, Netherlands"}}
Could you tell me if I need an umbrella for my walk today around the canals of Amsterdam, Netherlands?
[…]
{"CurrentApp":"Weather","Config":{"City": "Amsterdam, Netherlands"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "Cape Town, South Africa"}}
I’m packing for a business trip and I’m not sure what to bring. What’s the current weather in Cape Town, South Africa?
[…]
{"CurrentApp":"Weather","Config":{"City": "Cape Town, South Africa"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "Venice, Italy"}}
Just finished a book set in Venice, Italy, and it got me wondering about the weather there today. Is it as sunny as in the story?
[…]
{"CurrentApp":"Weather","Config":{"City": "Venice, Italy"}}
[…]
{"CurrentApp":"Weather","Config":{"City": "the Great Pyramids stand"}}
My friend in Sydney mentioned they were having quite the heatwave. I’d love to compare that to the current temperatures where the Great Pyramids stand. Could you give me a quick overview of the weather there?
[…]
{"CurrentApp":"Weather","Config":{"City": "the Great Pyramids stand"}}
$\times$
[…]
{"CurrentApp":"Calculator","Config":{"promptSequence": "24 / 6"}}
I’ve got 24 cupcakes, and I need to divide them evenly among my 6 friends. How many does each person get?
[…]
{"CurrentApp":"Calculator","Config":{"promptSequence": "24/6"}}
$\checkmark$
Table 2: Comparison of GPT-4 generated inputs with our custom multimodal LLM engine consisting of custom fine-tuned Google T5 & BERT-Large-Uncased-Whole-Word-Masking-Squad2 outputs. The model performs with extreme efficacy on descriptive and lengthy prompts. However, it can be observed to struggle with advanced mathematical prompts as it is strictly limited to segment extraction. Mathematical prompts can come in all sorts of wrangled manners, and the BERT model cannot handle expression extraction with logical implications. This is why the custom-tuned T5 proves to be so powerful, with all tasks.
6 Future Directions
To improve this framework and scale its potential, continuous iterations on both front-end and back-end engine components need to be considered.
6.0.1 Automation of Annotated UI Paradigm
An automated structure for generating classical UI components through textual descriptions would allow for a unified annotated system. If a developer’s textual prompts can generate UI components, they can be parsed to serve as annotations for the components as well. This would immediately make the UI compatible with the LMM engine’s requirements and would transform the way full-stack applications are developed.
6.0.2 LLM Engine Embellishments
Exploring configurations regarding transformer models that classify, extract and parse information in the LLM engine will improve accuracy and inference speed. The choice between delegating tasks to different models versus using a singular multi-task model for all tasks is open to be explored.
6.0.3 AutoGen
Expanding on the back-end embellishments, making use of Microsoft’s AutoGen (Wu et al. (2023)), Mistral’s Mixtral 8x7B mixture of experts model (Jiang et al. (2024)), or Ollama’s model hot-swapping to allow for multi-model agent setups may be the ultimate solution for delegating a task to the appropriate model rather than having a singular model handle all tasks. If an AutoGen setup can delegate a task to the best-suited LLM available in the engine, accuracy can be maximized across all tasks and inference times can be dynamic depending on the complexity of the task, rather than potentially sluggish times even for simpler tasks that don’t require larger models.
6.0.4 Concurrent Processing
Further scaling would call for large-scale concurrent processing of incoming user requests. Whilst our backend server is currently multi-threaded and can handle many inputs simultaneously, further scaling to incorporate asynchronous processing with expanded capabilities, along with more robust load balancing and distributed computing systems, would bulletproof our framework’s servers.

### v2

@@
 Abstract
-The recent meteoric advancements in large language models have showcased a remarkable capacity for logical reasoning and “comprehension”.
-These newfound capabilities have opened the door to a new generation of software, as has been made obvious through the innumerable ways they are being applied in the industry.
-This research focuses on harnessing and guiding the upgraded power of LLMs to construct a framework that can serve as an intermediary between a user and their user interface.
-By comprehending a user’s needs through a thorough analysis of natural textual inputs, an effectively crafted LLM engine can classify the most likely available application, identify the desired UI component and subsequently execute the user’s expected actions.
-This integration can evolve static UI systems into highly dynamic and adaptable solutions, introducing a new frontier of intelligent and responsive user experiences.
-Such a framework can fundamentally shift how users accomplish daily tasks, skyrocket efficiency, and greatly reduce cognitive load.
+The evolution of Large Language Models (LLMs) has showcased remarkable capacities for logical reasoning and natural language comprehension.
+These capabilities can be leveraged in solutions that semantically and textually model complex problems.
+In this paper, we present our efforts toward constructing a framework that can serve as an intermediary between a user and their user interface (UI), enabling dynamic and real-time interactions.
+We employ a system that stands upon textual semantic mappings of UI components, in the form of annotations.
+These mappings are stored, parsed, and scaled in a custom data structure, supplementary to an agent-based prompting backend engine.
+Employing textual semantic mappings allows each component to not only explain its role to the engine but also provide expectations.
+By comprehending the needs of both the user and the components, our LLM engine can classify the most appropriate application, extract relevant parameters, and subsequently execute precise predictions of the user’s expected actions.
+Such an integration evolves static user interfaces into highly dynamic and adaptable solutions, introducing a new frontier of intelligent and responsive user experiences.
+Keywords Large Language Models, Context-Aware Pervasive Systems, Semantic Modelling, Agent-Based Reasoning, Human-Computer Interaction
 1 Introduction
 The modern world relies on and is driven by software.
-Embedded systems, command-line interfaces and user interface (UI) software are present across systems all around the world.
-The ease of use coupled with their intuitive nature has allowed for UI systems to become a staple as a crucial tool in modern software and beyond.
-UI systems serve as a visually appealing packaging of function calls and event handlers, allowing for complex event pipelines and data flows to be abstracted by buttons, text fields, menus, etc.
-The evolutions made in large language models (LLMs) over the past year have exhibited true "cognitive" potential.
-This potent ability has unveiled innumerable new opportunities to revolutionize the way our contemporary software systems are expected to operate.
-In this paper, we explore our vision and progress toward developing a UI architectural paradigm which employs a multimodal engine powered by LLMs and state-of-the-art transformer models.
-This framework aims to abstract monotonous UI interactions with prompting mechanisms that serve as "cognitively aware", powering automated functional calling and data flow pipelines, which translate to full speech-based intelligence control over visual UI systems.
-Our approach posits a novel architecture and methodology for UI systems, incorporating logically sound LLM engines capable of interpreting a user’s desired action.
-This facilitates a synergistic interaction between the software and the user, enabling a more intuitive and user-centred experience, rather than requiring the user to conform to the software’s predefined structures.
-2 Motivations of the Research
-The catalyst for this research is grounded in the critical examination of current user interface systems, which, while omnipresent and essential, do not come without limitations.
+Embedded systems, command-line interfaces, web and desktop applications are present across systems all around the world.
+The predominant way in which we interact with software is through user interfaces.
+Their ease of use coupled with their intuitive nature has allowed for user interfaces to become an indispensable tool across contemporary software.
+These systems serve as a visually appealing packaging of function calls and event handlers, allowing complex event pipelines and data flows to be abstracted by buttons, text fields, menus, and so on.
+The catalyst for our research is grounded in the critical examination of current user interface systems, which, while omnipresent and essential, do not come without limitations.
 Traditional implementations of UIs often require the user to adhere to the methods and system constraints established by the developer and their design.
 This means that all UI components in an application have specific triggers and their invocation is limited to a strict sequence of user inputs within the applications’ available features.
-Our paradigm would relieve the user of such restricted, stringent and static rule sets when utilizing software, fundamentally changing the landscape of how applications are developed, used and integrated across all sectors.
+The substantial advancements made in the space of large language models, especially over the past few years, have exhibited a true “cognitive" potential.
+This potent ability has unveiled innumerable new opportunities to revolutionize the way our software systems are expected to operate.
+It became clear to us that an end-to-end architecture employing textual semantic modelling of contemporary user interfaces, as well as leveraging familiar user prompting behaviour, could be crafted.
+Within our research, we explore our vision and progress toward developing such a UI architectural paradigm, which benefits from the integration of a custom multimodal engine powered by LLMs and state-of-the-art transformer models.
+Our proposed framework aims to textually model the semantics of UI components, giving them the ability to communicate effectively with an LLM-driven agent environment.
+The automated generation of metadescriptors and annotations modelling each UI component allows us to form a data structure that streamlines agent-based parsing and entity extraction.
+Through this, we facilitate a synergistic interaction between the software and the user, enabling a more intuitive and user-centred experience, rather than requiring the user to conform to the software’s predefined structures.
+Thus, our framework aids in abstracting monotonous UI interactions with prompting mechanisms that serve as “cognitively aware", powering automated function calling and data flow pipelines.
+This translates to fully speech-based and intelligent control over visual user interfaces.
+The benefits of our paradigm and its extensions are vast.
+Firstly, it would relieve the user of restricted, stringent and static rulesets when utilizing software.
+This would fundamentally change the landscape of how applications are developed, used and integrated across all sectors.
 It would allow for real-time intelligent and efficient control over complex tasks, enabling users to achieve substantially more, in significantly less time.
-Our research aims to foster this innovation by bridging the gap that exists between the research and practical integration of LLMs into UI systems.
-Through development and implementation, valuable empirical insights allow for a demonstration of how LLMs can not only integrate with UI systems but also revolutionize them.
-From enhancing accessibility for users with disabilities to streamlining complex workflows in enterprise systems, the implications are vast and worthy of exploration.
-This paper seeks to motivate and propagate a new wave of research and development in this direction, to realize the full potential of LLMs in creating more adaptive, intuitive, and efficient UI systems.
-This combination of deep theoretical exploration and practical application is vital for the next generation of software development.
+They would be capable of multitasking through speech and executing complicated operations with real-time UI responses.
+Secondly, such a framework would further democratize software by drastically reducing learning curves, specifically for intricate software.
+A user would be able to provide their intended actions through speech and textual input and delegate heavy interactions to the framework, letting it reason with the UI to get the job done.
+Another major benefit of this framework is the streamlining of application development.
+User interfaces would be simplified substantially, hosting only the most necessary components.
+Furthermore, developers can generate annotations efficiently or automatically, allowing them to integrate functionality with far fewer input UI components.
+The paper is organized in the following manner.
+In section 2, we showcase related works that introduce what has already been achieved in related domains, as well as where we fill gaps that are present.
+In section 3, relevant challenges in the development of our framework and how they were overcome are thoroughly explained.
+Section 4 is where our framework is formally proposed in detail.
+Section 5 discusses evaluations of the framework’s accuracy and precision, along with showcasing comparisons between the integration of various LLM models.
+Finally, sections 6 and 7 discuss future directions, final thoughts, and conclude the paper.
+2 Related Work
+The interest in applying large language models to user-centred domains is only natural, given that the greatest strength of these models is their semantic comprehension and reasoning abilities.
+Many research endeavours are focused on using LLMs in the context of human-computer interaction, crafting solutions for increased adaptation to user actions and behaviour.
+This includes training LLMs to better incorporate prompts from users through guided training pipelines which facilitate accurate execution of user commands [1], contextualization embeddings which grant more personalized responses [2], as well as implementing low-code interaction frameworks offering more intuitive control mechanisms over LLM instructing [3].
+These works have pushed the capabilities of LLMs forward and opened up more powerful and efficient solutions for user-focused systems.
+Other efforts are being made in making LLMs increasingly flexible and versatile tools for human-computer interaction, pushing for “journey" and context-aware interactions through few-shot learning methods [4], analysis of user-interest journeys, and desires [5].
+Training language models to teach themselves how to use tools in a self-supervised fashion and enabling zero-shot performance through demonstrations of API tools has also broadened the horizons of how LLMs can be integrated into existing pipelines [6].
+Many works, similar to those mentioned, are geared towards direct interactions with LLMs, aiming to alleviate reliance on or bypass traditional UIs in favour of enhanced one-on-one sessions between the user and LLM systems.
+This can significantly improve efficiency and enable more complex operations.
+However, we find that there is a lack of focus on adapting and evolving contemporary user interfaces.
+While many system implementations would benefit from the dismissal of UIs, the focus on adding intelligence to their operations is crucial for the next generation of software.
+Users need to be able to visually observe their data as well as the UI’s responses to their commands.
+They must be able to easily provide follow-up instructions with confidence that the framework will adapt in real time, equipping them with the ability to complete operations, simple or complex, with equal convenience.
+Eliminating learning curves or time-intensive tasks through intelligent speech control is invaluable.
+These factors are what led us towards the research and creation of a two-component framework that revolutionizes both ends of the system.
+A paradigm that changes the way we develop applications, harnessing the power of LLMs to comprehend user prompts and permit back-and-forth agent-based reasoning with our dynamic annotation tree that semantically models the nature of all available UI components.
+This approach not only facilitates task efficiency, customization, and user-adaptive behaviour, but also simplifies user interface development and consumption, allowing applications to deliver valuable context, user control, and accessibility.
+We aim to leverage this architecture to make technology more approachable and usable for a broader audience.
 3 Challenges
@@
 An effective fusion of LLMs and event-driven UIs is a complex task with many critical factors of concern.
-All classifying, processing, parsing and model inference must be achieved in real-time, the architecture must be capable of scaling and integrating unseen sub-applications in zero-to-few shots, and the framework’s interpretation accuracy of ambiguous user inputs must be top-notch.
-The greatest challenges encountered during our research and implementation of this framework are as follows:
+All classification, processing, parsing, and model inference must be achieved in real-time and the architecture must be capable of scaling and integrating unseen sub-applications in zero-to-few shots.
+Additionally, the framework’s interpretation accuracy of ambiguous user inputs must be top-notch.
+The greatest challenges encountered during our research and implementation of this framework are discussed in this section.
 3.1 Integration Complexity
-The underlying challenge across the entire research was effectively fusing LLMs with event-driven UIs.
-Developing an effective framework capable of being scaled with increasing application counts and fostering potential for implementation as an operating system was always a priority.
+The underlying challenge across the entire research endeavour was effectively fusing LLMs with event-driven UIs.
+Developing an effective framework capable of being scaled with increasing application counts, as well as fostering potential for implementation as an operating system was always a priority.
 A paradigm as dynamic as the software it intends to power was vital.
 Implementing a brute-force hard-coded solution would not serve a great purpose or carry the potential for scalability and impact required for next-generation software.
-This is why we took our time and carefully crafted each component of this framework with future scaling in mind, leaving room for future advanced integrations such as multi-agent systems as well as more powerful LLM model configurations.
+This is why we took our time and carefully crafted each component of this framework with scaling in mind, leaving room for future advanced integrations, such as more powerful LLM model configurations and automated agent fine-tuning.
 3.2 Annotation Data Structure Selection & Traversal
-Another critical point of concern was the establishment of a data structure suitable and most effective for representing the overarching system, applications, as well as all UI components per application.
+Another critical concern was selecting a data structure to effectively represent the system, its applications, and all UI components.
 Through our research, a tree structure made for the most accurate representation of the overall system, whilst also allowing for efficient traversal and information retrieval, vital for real-time responses further on.
@@
 Rather than mapping an input to an application and its UI components through an unstructured or linear search, an organized level-by-level search eliminates redundant checks.
-Each user input is compared to a tree, which stores textual naming, descriptions, expectations, and examples of each application and its UI components.
+Each user input is compared to a tree, which stores textual naming, descriptions, expectations, and examples of every application and its respective UI components.
 Our use of a top-down manner comparison between the tree structure and user input ensures that only children of relevant nodes are searched, preserving our engine’s real-time processing goals.
-Listing 1: An example of how an application’s component metadata is stored in the program’s node structure
-Extract the location: "I am Brian O’Connor residing at Apartment 5A, 654 Peachtree Street in New Town and my email address is brian.oconnor@mailservice.net." || "Address": "Apartment 5A, 654 Peachtree Street in New Town"’,
+Listing 1: An example of how an application’s component metadata is stored as a node in the annotation tree data structure
+Extract the name, location, and email: "I am Brian O’Connor residing at Apartment 5A, 654 Peachtree Street in New Town and my email address is brian.oconnor@mailservice.net." || "Name": "Brian O’Connor","Address": "Apartment 5A, 654 Peachtree Street, New Town", "Email": "brian.oconnor@mailservice.net"’,
 ‘Extract the arithmetic expression: "If you have \$50 and spend \$25, how much money do you have left?" || "\$50 - \$25"’,
@@
 3.3 Model Selection & Training
-For our framework, an engine that is capable of high-level logical reasoning, understanding and general task capability was paramount.
+For our framework, an engine that is capable of high-level logical reasoning, comprehension, and general task capability was paramount.
 A multimodal pipeline was found to be most efficient in allowing for the balancing of inference times between models with parameter sizes appropriate for the task, as opposed to a single overkill model wasting processing time on smaller classification jobs.
-Tasks that could be accurately and fastest completed by smaller models were assigned a certain model, while more complex tasks such as input comprehension, parameter extraction and parsing were given their own appropriate models.
-For each task, many models were researched, to balance accuracy and speed.
-We found ROBERTA models (Liu et al.
-(2019)) fine-tuned on datasets such as SQUAD2 (Rajpurkar et al.
-(2018)) were efficient at the task.
-For comprehension and parsing upon our application library, we found that even powerful and large models such as Google’s XL T5 (Raffel et al.
-(2023)) pre-trained to perform well with general multi-task capabilities simply cannot achieve the desired accuracy across our simple application library.
+Tasks that could be accurately and fastest completed by smaller models were assigned models accordingly, while more complex tasks such as input comprehension, parameter extraction and parsing were given their own appropriate models.
+For each task, many models were researched with the goal to balance accuracy and speed.
+We found ROBERTA models [7] fine-tuned on datasets such as SQUAD2 [8] were efficient at entity extraction tasks in terms of speed, yet lacked precision when challenged with convoluted extractions.
+For such comprehension and parsing, we found that even powerful and large models such as Google’s XL T5 [9] pre-trained to perform well with general multi-task capabilities simply cannot achieve the desired accuracy across our application library.
 The clear answer would be to employ larger and more powerful models, however, inference times and computational resources quickly become a concern.
 Custom fine-tuning was key and opened a pathway for us to research how to achieve generalization well enough to output consistently formatted and precise results.
+See Table 1.
 While training models on custom datasets relevant to each application in our library is effective, a scalable and more promising approach would be achieving few/zero-shot capabilities.
@@
 Table 1: Model Benchmarks
-T5
-Custom Fine-Tune
-T5
-Google MTL
-BERT
-SQuAD2
-ELECTRA SQuAD2
+T5 Custom Fine-Tune T5 Google MTL BERT SQuAD2 ELECTRA SQuAD2
 Weather 95% 36% 96.8% 76%
@@
 3.4 Training Data Generation & Wrangling
-Creating custom datasets with realistic, convoluted and vague prompts reflective of real-life scenarios was tough.
-We found no available datasets with the level nature of prompts required to be effective for our training.
+Creating custom datasets with realistic, convoluted and vague prompts reflective of real-life scenarios was challenging.
+We found no available datasets with the nature of prompts required to be effective for our training.
 Custom datasets were the best option and were a significant task.
-This is where generation with top-of-the-line LLMs such as ChatGPT 4 served immense value.
-Many competing models were tested, however, the adherence to prompt engineering, logical understanding of the task, as well as the depth of versatility GPT 4 (OpenAI (2024)) was able to provide, we found to be unmatched.
+This is where generation with top-of-the-line LLMs such as ChatGPT-4 served immense value.
+Many competing models were tested, however, the adherence to prompt engineering, logical understanding of the task, as well as the depth of versatility GPT-4 [10] was able to provide, were unmatched in our findings.
 OpenAI’s API was used for mass data generation and any formatting inconsistencies were fixed by hand.
-These custom datasets were used to fine-tune LLMs such as varying sizes of Google T5 on various tasks.
-4 Proposed LM-UI Integration Framework
-With the aforementioned motivations and challenges, we present the LM-UI framework.
-The integration of a hybrid design that merges an annotated front-end component system with a multimodal classification and logical extraction engine, allows us to achieve a strong logical next-generation user experience (UX).
-Our framework features a UI paradigm that dispatches its component event handler states to a central Redux store.
-This store then uses reducers to communicate with the LLM engine, appropriately allowing updates to the corresponding front-end state.
-This framework allows UI software to interpret and predict a user’s target action, even when provided vague or convoluted prompting.
-4.1 Modelling UI Controls
-Our approach to fusing LLMs with UI hinges on the modelling of UI components as adaptive entities capable of dynamically responding to natural language inputs from the user.
-At the core of this integration is a re-imagined method to define UI elements such as buttons, text fields, navigation links, etc.
-These components are incorporated to be part of an event-driven environment where their behaviours and states are modifiable based on user interactions conveyed through natural language prompts.
-4.1.1 Meta Description of UI
-Meta descriptions of each UI component present in an application are critical to accuracy in our framework.
-The use of textual metadata providing detailed specifications of how each UI component fits conceptually within the structure of the application, what its configurable parameters are and examples of potential use cases, prime the application for communication of its expectations and needs with the LLM engine down the road.
-4.1.2 Descriptive Component Modeling
-Each application is treated as a "node", with all its UI elements being "sub-nodes".
-These nodes serve as a way to represent the structure of an application.
-All nodes have textual descriptions outlining their purpose in the application, while sub-nodes have descriptions or example questions related to how they may be prompted for extraction during inference.
-These details allow our LLM engine to understand and have reference knowledge as to what each application does and wants.
-Listing 2: An example of how an application’s component metadata is stored in the program’s node structure
-NODES = {
-    "Weather": {
+These custom datasets were used to fine-tune LLM models, such as varying sizes of Google T5 on various tasks.
+4 Proposed LM-UI Framework Architecture & Design
+With the aforementioned motivations and challenges, we present the Language Model-User Interface framework.
+Through the fusion of an annotated front-end component system and a multimodal extraction engine specialized in classification and logical entity recognition, we are able to lay the groundwork for the next generation of user experiences.
+The framework employs an end-to-end custom architecture that hinges upon the integration of a specialized tree data structure which effectively manages annotations and metadescriptions of UI components.
+The creation, indexing, and merging of this data structure is what allows us to employ LLM agents capable of real-time interpretation and prediction regarding a user’s intended actions.
+Through this pipeline, user interfaces can be significantly simplified, whilst becoming increasingly dynamic and efficient.
+Figure 2: Two-Component Framework
+4.1 Describing Applications as Events & States
+The first aspect of our architecture asks that the applications in our framework be modelled upon the concept of events and states.
+Each UI element in an application has a state for all its configurable parameters.
+Within each application’s root component, its states are observed for any changes.
+When states are updated in response to a user event, the component is immediately re-rendered to reflect the new states.
+When the master JSON object holding a current application indicator along with the states of its respective parameters is updated, the new states are dispatched to the central store.
+Each interaction with the UI components of an application is an event that is reflected by the Redux states and store.
+4.1.1 Central Store & Reducers
+When a user’s textual prompt is received, it is sent to the LLM engine for processing.
+Once the engine returns the parsed result to the root front-end component, it triggers action events.
+These events dispatch the new states to reducers which are functions that take the previous state and an action as arguments and calculate a new state.
+They are responsible for handling transitions from one state to another within the application, ensuring state changes are predictable and transparent.
+The Redux central store is then updated to reflect the new state.
+When this store is updated, useSelector hooks and connect functions detect the state change and immediately update the UI.
+Figure 3: Application State Flow
+In the context of our framework, when an action is dispatched to the Redux store, it is forwarded to the corresponding reducer.
+The reducer, being a pure function, takes the current state and the dispatched action as its arguments.
+It then calculates and returns a new state without modifying the old state.
+The Redux store updates its internal state with this new state returned by the reducer.
+This updated state then propagates through the application, automatically triggering updates in the UI components that are dependent on this part of the state.
+This process facilitates a smooth, unidirectional data flow, ensuring that UI components can remain stateless, which promotes their reusability and testability.
+4.1.2 Case Study: Application/Task Library
+The implementation of this framework serves as a type of “operating system” that functions over a library of applications.
+Thus far, we have three simple applications in the library: An account sign-up form, a live weather application, and a calculator.
+These applications help showcase and validate the core functionality and flexibility of our architecture.
+It is easy to see how this can be scaled up to include various essential applications that are commonly found on smartphones, tablets, smartwatches, etc.
+Once we can establish and validate the framework’s potential and ability, more complex and intricate UI applications can be tackled with the appropriate merging and implementation schemes.
+4.2 Modelling User Interface Components
+Our approach to fusing LLMs with UI stands firmly upon the modelling of UI components as adaptive entities capable of dynamically responding to natural language inputs from the user.
+At the core of this model is a re-imagined method to define UI elements such as buttons, text fields, navigation links, etc.
+These components are thoroughly described and textually modelled in our annotation tree data structure.
+As a result, their integration and flexibility are streamlined in an event-driven environment where behaviours and states are modifiable based on user interactions conveyed through natural language prompts.
+Figure 4: Each application is semantically mapped into the tree structure. Each node holds detailed annotations
+4.2.1 Annotation Tree & Meta Descriptions
+The data structure storing meta descriptions of each UI component in an application is the key principle behind the proposed framework.
+Our requirement of this textual modelling is the crucial link bridging the gap between contemporary user interfaces and the emergence of highly advanced LLMs.
+These annotations provide detailed specifications of how each UI component fits conceptually within the structure of the application, what its configurable parameters are, and examples of potential use cases.
+This information primes applications for communication with the back-end engine’s agents further down the pipeline.
+Let us denote our application library system as $S$, where $S$ contains all elements representing applications and their components:
+$S=\{A,W,C\}$
+where A, W, and C denote the sets for the Account Form, Weather, and Calculator applications from our application library, respectively. Let each nested application set be denoted as:
+$A=\{a_{1},a_{2},a_{3}...\}\textit{ for Account Form components}$
+$W=\{w_{1},w_{2},w_{3}...\}\textit{ for Weather components}$
+$C=\{c_{1},c_{2},c_{3}...\}\textit{ for Calculator components}$
+where each contained element represents a unique node modelling a UI component.
+Each node stores annotations describing the component’s role in the application, along with the descriptors of what parameters it manipulates.
+See Listing 1.2
+Listing 2: An example of how an application’s component metadata is stored as a node in the annotation tree data structure
+a1: {
 "AppName": "Weather",
+"Component Name": "Location Field"
 "Description": "A sub-node of Weather. Represents the point where location or query details are input for weather-related inquiries.
 "Parameters": { "City": "What is the location?"}
-4.2 Describing Applications as Events & States
-To harmonize our front-end applications with our descriptive component modelling, the applications in our framework are modelled upon the concept of events and states.
-Each UI element has a state for all its configurable parameters.
-Within each application’s parent component, its states are observed for any changes by React useEffect hooks.
-When states are updated in response to a user event, the useEffect hook immediately updates the UI to reflect the new states.
-When the master JSON object that holds the current application indicator along with the states of its respective parameters is updated, it is dispatched to the central store.
-Each interaction with UI components of an application is an event that is reflected by the redux states and store.
-4.2.1 Central Store & Reducers
-The Redux store serves as the central point for handling each one of our applications’ states.
-When a user’s textual prompt is received, it is sent to the LLM engine to be processed.
-Once the engine returns the parsed result to the root front-end component, it triggers action events.
-These events dispatch the new states to the Redux central store where reducers are employed.
-The reducers are functions that take the previous state and an action as arguments and return a new state.
-They are responsible for handling transitions from one state to another state within the application, ensuring state changes are predictable and transparent.
-In the context of our framework, when an action is dispatched to the Redux store, the corresponding reducer processes the action, updates the state based on the action type and payload, and returns the updated state.
-This new state then propagates through the application, triggering UI updates where necessary.
-The use of reducers thus facilitates a smooth, unidirectional data flow and this separation of concerns ensures that the UI components are kept as stateless as possible, promoting reusability and testability.
-4.2.2 Case Study: Application/Task Library
-The implementation of this framework serves as a sort of “operating system” that functions over a library of applications.
-Thus far we have three simple applications in the library: An account sign-up form, a live weather application, and a Calculator.
-These applications help showcase and validate the core functionality and flexibility of our framework.
-It is easy to see how this can be scaled up to include various essential applications that would be commonly found on smartphones, tablets, smartwatches, etc.
-Once we can establish and validate the framework’s potential and ability, more complex and intricate UI applications can be tackled with the appropriate implementation schemes.
-4.3 Natural Language Parser
-Our LLM engine stands firmly on the foundation established by the preceding components and its ability is solely reliant on the groundwork we laid out.
-When natural language prompts from the user are sent to the engine, it traverses the following pipeline:
-Figure 2: Two-Component Framework
-4.3.1 Tokenization
+4.3 Parsing & Mapping User Input
+The user provides input describing their target intention through the system’s textual prompt input field or through a speech button which triggers a recording of the user’s spoken prompt.
+Once received and converted to text through speech recognition models, if speech is used, a prompt is sent to the back end where it is processed by an agent-based prompting mechanism.
+The processing pipeline carries out appropriate encoding, tokenization, and finally mapping via classification.
+Tokenization
   1. 1.
-    An input sentence is split into tokens. BERT uses WordPiece tokenization
+    An input sentence is split into tokens, adhering to an agent’s specific LLM encoding requirements. We will cover BERT and its use of WordPiece tokenization for this example
   2. 2.
@@
 […]
-4.3.2 Token to ID Conversion
+Encoding
   1. 1.
-    Each token is then mapped to an ID using BERT’s vocabulary. This can be represented as:
-    $\text{ID}_{i}=\text{Vocab}(\text{"Token}_{i}\text{")},\text{ for }i=1,2,...,N$
+    Each token is then mapped to an ID using BERT’s vocabulary. This can be represented as
+    $\text{ID}_{i}=\text{Vocab}(\text{``Token}_{i}\text{")},\text{ for }i=1,2,...,N$
     where ‘Vocab‘ is a function that returns the ID of a token in the BERT vocabulary.
-4.3.3 Segment and Positional Embeddings
+  2. 2.
+    The total embedding for a token, $E_{token}$, includes word, segment, and positional embeddings.
+4.3.1 Mapping via Classification
+To classify user input, we encode the input and node representations:
+  • For user input $u_{i}:v_{u_{i}}=encode(tokenize(u_{i}))$
+  • For a node $n_{j}$ representing an application or component: $v_{n_{j}}=encode(tokenize(n_{j}))$
+Let U denote the user’s encoded input:
+$U=\{u_{1},u_{2},u_{3}...\}$
+The system then maps user input to system nodes using an interpretation function that performs classification, $f:U\longrightarrow S$, where:
+$f(U_{i})=n_{k}$, where $n_{k}$ is chosen such that:
+$n_{k}=argmax(cosine\_similarity(\vec{v}_{u_{i}},\vec{v}_{n_{j}}))$
+This process classifies the user’s input to the most relevant application component in the annotation tree.
+To handle complex inputs that involve specific actions or components, we extend the function $f$ to further map inputs to nested sets (components) within A, W, or C:
+$f(u_{i})=a_{j}\in A\textit{ or }f(u_{i})=w_{j}\in W\textit{ or }f(u_{i})=c_{j}\in C$
+This implies that the function not only identifies the relevant application but also the specific action or component within that application.
+4.3.2 Entity Extraction and Parameterization via LLM Inference
+At this level, the custom fine-tuned LLM agents with specialized ability such as Google’s T5 model trained on specific tasks from our library in a teacher-student environment, enable accurate parsing of complex user inputs into structured data that the system can act upon.
+This process is crucial for interpreting the specific details contained within the user’s input and mapping them to actionable UI component parameters in our system.
+Figure 5: Entity Extraction & Paramaterization
+For inputs that require extracting parameters or entities, we define a transformation $g$: $U$$P$, where P is a set of parameters extracted from the inputs.
+This can be used alongside $f$ to not only map an input to an application component but also extract the necessary information to perform the action:
+$\textit{g}(\textit{u}_{\textit{i}})=\{\textit{p}1,\textit{p}2,…\}$
+where $u_{i}$ is the encoded user input, and $\{p_{1},p_{2},...\}$ are the extracted parameters.
+These parameters are then used to dynamically adapt the system’s response to the user’s intent.
+This extraction algorithm functions as a decision-making tool with the use of natural language processing techniques.
+Its objective is to analyze textual input, categorize it based on predefined criteria, and generate a JSON parameter extraction object to be relayed to the front end and dispatched to the Redux central store.
+4.3.3 Action Execution Based on Mapping
+Finally, we use the outcomes of f and $g$ to create a final JSON object and execute actions within the system.
+This involves sending the final JSON object to the front end where the Redux central store invokes specific methods associated with the identified components, passing the extracted parameters and dynamically updating the UI.
+Let the following represent the introduction of relevant methods:
+$execute(f(u_{i}),g(u_{i}))$
+4.3.4 Example:
+Assume a user input $u_{i}$= “I have a flight to France, what’s it like out there?".
+The interpretation function f maps this input to the weather application’s “location" field component $w_{location}\in W$, and $g$ extracts “France" as a parameter.
+The system then executes the location-based action for “France" based on this mapping and parameter extraction.
+The defined detailed approach allows the system to dynamically interpret user inputs by understanding their intent and mapping them to the most relevant application and components in the application library system.
+It ensures that user inputs are handled efficiently, navigating through the hierarchical structure of the application library based on semantic similarity, thereby improving user experience through accurate identification and interpretation of a user’s needs.
+4.4 Operational Workflow
   1. 1.
-    For BERT, segment embeddings (to differentiate between different sentences) and positional embeddings (to give information about the position of the token in the sequence) are also added
+    Input Analysis The algorithm initially analyzes the input to determine its nature and relevance to each UI component child node under the classified application.
+This categorization is crucial in deciding the subsequent processing path.
+    1. (a)
+      Top-Down, First-Match Traversal The system starts the mapping at the root of the application tree.
+It uses our defined interpretation function f to select the child node with the highest cosine similarity to the user input.
+The process repeats until a leaf node is reached, moving to children nodes only after the parent node is classified.
+    2. (b)
+      Early Stopping for Efficiency & Error Mitigation A threshold is in place for cosine similarity.
+If the similarity between the user input and the best-matching node at any level is below a predefined threshold, the search is halted, and the system can proceed with action execution based on the current best match, or prompt the user to try again.
+This optimizes the search process, ensures responsiveness and mitigates errors that can result from forced mapping.
   2. 2.
-    The total embedding for a token is then:
-    $E_{\text{token}}=E_{\text{word}}(ID)+E_{\text{segment}}+E_{\text{position}}$
+    Conditional Logic Based on the analysis, the algorithm employs conditional logic to deploy appropriate agents in the extraction mechanism.
+Parent nodes of classified applications are checked, to guarantee the most suitable model is utilized.
+This selection process is pivotal to the algorithm’s adaptability and accuracy in parameter extraction.
   3. 3.
-    Here, $E_{\text{word}}(ID)$ is the word embedding corresponding to the token ID, $E_{\text{segment}}$ is the segment embedding, and $E_{\text{position}}$ is the positional embedding.
-    Cosine $\text{Cosine Similarity}(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$
-The cosine similarity between a user’s input and an annotated node will determine the likelihood that the user desired that node.
-Whichever node has the highest similarity with the user’s input will end up being the classified application, and the engine will traverse its children nodes as a result.
-This type of Best-First Search mitigates redundant comparisons through all leaf nodes from incorrect applications, in turn significantly reducing the chance for incorrect matching with incorrect UI components.
-Parameter Extraction
-Proceeding down the pipeline is the core function of the engine.
-At this level, a powerful LLM with substantial generalization and multitasking ability is used to compare the annotated children nodes of the classified application and compare each to the user’s input.
-Through prompt engineering, an LLM such as Google’s T5 model fine-tuned on large text corpora for accurate NLP multitasking abilities, can determine which part of the user’s input corresponds to the current child node being checked.
-This extraction algorithm functions as a decision-making tool with the use of natural language processing techniques.
-Its objective is to analyze textual input, categorize it based on predefined criteria, and generate a JSON parameter extraction object, to be relayed to the front end and dispatched to the Redux central store.
+    Response Generation Depending on the categorization:
+    • If the input falls into a specific category (e.g.
+calculations), the algorithm engages a dedicated module designed for that category, employing specialized deep-learning models, such as our custom fine-tuned Google’s T5
+    • For other types of input, the algorithm utilizes a different set of models tailored for general language understanding and response formulation
+    • This delegation of tasks makes it so that we use the most appropriate model for the task, employing larger more capable models for more demanding tasks such as complex arithmetic extraction, while making use of suitable smaller models for simplistic tasks.
+This optimization ensures valuable inference time from larger models is not wasted on tasks that don’t require it, saving resources along with crucial response time.
+  4. 4.
+    JSON Output Formatting Finally, the parameters extracted by the model are formatted into a structured JSON object, optimized and suitable for further processing or for display to the end-user once sent back to the client.
+See Listing 1.3.
 Listing 3: Extracted parameters parsed into a JSON object
 {"CurrentApp":"AccountForm","Config":{"Name": "Connor Syle", "Address": "34 Coronation Street", "Email": "connor32@outlook.com"}}
-4.4 Operational Workflow
-  1. 1.
-    Input Analysis: The algorithm initially analyzes the input to determine its nature and relevance to each UI component child node under the classified application.
-This categorization is crucial in deciding the subsequent processing path
-  2. 2.
-    Conditional Logic: Based on the analysis, the algorithm employs conditional logic to select an appropriate extraction mechanism.
-Parent nodes of classified applications are checked, to guarantee the most suitable model is employed.
-This selection process is pivotal to the algorithm’s adaptability and accuracy in parameter extraction.
-  3. 3.
-    Response Generation: Depending on the categorization:
-    • If the input falls into a specific category (e.g., calculations), the algorithm engages a dedicated module designed for that category, employing specialized deep-learning models, such as our custom fine-tuned Google’s T5
-    • For other types of input, the algorithm utilizes a different set of models tailored for general language understanding and response formulation
-    • This delegation of tasks ensures we use the appropriate model for the task, employing larger more capable models for more demanding tasks such as complex arithmetic extraction, while making use of suitable smaller models for simplistic tasks.
-This optimization ensures inference time is not wasted on tasks that don’t require it, saving resources along with crucial response time.
-  4. 4.
-    JSON Output Formatting: Finally, the parameters extracted by the model are formatted into a structured JSON object, optimized and suitable for further processing or display to the end-user once sent back to the client.
-See Listing 3
 This pipeline employed by our LLM engine allows for complex comprehension of textual input from a user, by accurately classifying and extracting parameters that the target application requires.
-This engine is able to streamline the connection between the user and the software by taking the intentions and requirements of both into account, utilizing its logical reasoning to complete the task as it sees fit.
-This is a stepping stone towards the right direction in the next generation of software.
+Our engine is able to streamline the connection between the user and the software by accounting for the intentions and requirements of both, utilizing its logical reasoning to complete the task as it sees fit.
+This is a stepping stone in the right direction for the next generation of software.
 5 Evaluation
@@
 The engine must correctly classify which task the user’s prompt corresponds to, extract, parse, and output the expected parameters in the required JSON format.
+{adjustwidth}
+[…]
+Table 2: Comparison of GPT-4 generated inputs with our custom multimodal LLM engine consisting of custom fine-tuned Google T5 & BERT-Large-Uncased-Whole-Word-Masking-Squad2 outputs.
+The model performs with extreme efficacy on descriptive and lengthy prompts.
 5.1 Required Tasks
@@
 This classification task is extremely reliable and the chance of failure is observed to be negligible.
-The task that is critical to evaluate is the parameter extraction once the input’s corresponding task has been established.
+The critical task to evaluate is the parameter extraction once the input’s corresponding application has been established.
 Extensive experimentation and trials were conducted with plentiful transformer models from the Hugging Face hub.
@@
 5.2 Model Selection
-For this evaluation, we pit the most effective models found during initial testing, with available resources for training, hosting, deployment and inference in mind)—the T5 Text2Text Generator models along with BERT and ELECTRA.
+For this evaluation, we pit the most effective models found during initial testing, with available resources for training, hosting, deployment and inference in mind—the T5 Text2Text Generator models along with BERT and ELECTRA.
 The BERT and ELECTRA models were already fine-tuned on the SQuAD 2.0 dataset, which is a popular dataset used in fine-tuning models for logical reasoning and question-answering.
-On the other hand, the T5 model base and multi-task, trained by Google were also chosen due to their flexibility, scalability, and efficacy in abstraction and multi-task situations.
+On the other hand, the T5 model base and multi-task, trained by Google, were also chosen due to their flexibility, scalability, and efficacy in abstraction and multi-task situations.
 Figure 6: Accuracy comparison between models engineered for our application’s tasks.
 Each model was trained on logical and information extraction tasks, and further carefully prompt engineered to achieve the best performance possible for each task.
-Text2Text Generation Models were fine-tuned on multiple datasets to multitask excellence.
-The T5 custom model was specifically trained on specialized custom datasets relevant to our application’s tasks.
 5.3 Observation
-We can see from the following benchmarks in Figure 3 that the custom-trained T5 model does exceptionally well across all tasks, performing the best in three out of four tasks, with the last one being a close tie.
+We can see from the benchmarks in Figure 6 that the T5 model trained on specialized custom datasets relevant to specific application tasks does exceptionally well across all tasks.
+It performs the best in three out of four tasks, with the last one being a close tie.
 The BERT model also performs very well but struggles with the logical capabilities required for the advanced calculator prompts.
-Google’s multitask-trained model falls extremely short in more logically involved tasks while the ELECTRA model does poorly overall.
-Further training of the T5 model with more diverse and better quality data would allow it to be the most versatile and powerful model for our project.
-GPT Baseline VS Custom Engine Prompt Evaluation
-Truth Intention GPT-4 Generated Inputs Multimodal Parser Output Eval
-[…]
-{"CurrentApp":"AccountForm","Config":{"Name": "Alex P. Globetrotter", "Address": "7890 Nomad Path, Backpack City", "Email": "wanderlust.world@travelbug.com"}}
-Oh, before I forget, jot down my email, it’s wanderlust.world@travelbug.com, and people shout out to me as Alex P.
-Globetrotter when they see me.
-The address?
-It’s 7890 Nomad Path, Backpack City.
-The mailman has it tough, I tell you!
-[…]
-{"CurrentApp":"AccountForm","Config":{"Name": "Alex P. Globetrotter", "Address": "7890 Nomad Path, Backpack City", "Email": "wanderlust.world@travelbug.com"}}
-[…]
-{"CurrentApp":"AccountForm","Config":{"Name": "Alex J. Turner", "Address": "768 Rolling Rock Street", "Email": "alex.turns@rocknmail.com"}}
-I’m registered under the name Alex J. Turner, but everyone sends their regards to my place at 768 Rolling Rock Street, and for a quicker response, they hit me up at alex.turns@rocknmail.com.
-[…]
-{"CurrentApp":"AccountForm","Config":{"Name": "Alex J. Turner", "Address": "768 Rolling Rock Street", "Email": "alex.turns@rocknmail.com"}}
-[…]
-{"CurrentApp":"Weather","Config":{"City": "Amsterdam, Netherlands"}}
-Could you tell me if I need an umbrella for my walk today around the canals of Amsterdam, Netherlands?
-[…]
-{"CurrentApp":"Weather","Config":{"City": "Amsterdam, Netherlands"}}
-[…]
-{"CurrentApp":"Weather","Config":{"City": "Cape Town, South Africa"}}
-I’m packing for a business trip and I’m not sure what to bring. What’s the current weather in Cape Town, South Africa?
-[…]
-{"CurrentApp":"Weather","Config":{"City": "Cape Town, South Africa"}}
-[…]
-{"CurrentApp":"Weather","Config":{"City": "Venice, Italy"}}
-Just finished a book set in Venice, Italy, and it got me wondering about the weather there today. Is it as sunny as in the story?
-[…]
-{"CurrentApp":"Weather","Config":{"City": "Venice, Italy"}}
-[…]
-{"CurrentApp":"Weather","Config":{"City": "the Great Pyramids stand"}}
-My friend in Sydney mentioned they were having quite the heatwave.
-I’d love to compare that to the current temperatures where the Great Pyramids stand.
-Could you give me a quick overview of the weather there?
-[…]
-{"CurrentApp":"Weather","Config":{"City": "the Great Pyramids stand"}}
-$\times$
-[…]
-{"CurrentApp":"Calculator","Config":{"promptSequence": "24 / 6"}}
-I’ve got 24 cupcakes, and I need to divide them evenly among my 6 friends. How many does each person get?
-[…]
-{"CurrentApp":"Calculator","Config":{"promptSequence": "24/6"}}
-$\checkmark$
-Table 2: Comparison of GPT-4 generated inputs with our custom multimodal LLM engine consisting of custom fine-tuned Google T5 & BERT-Large-Uncased-Whole-Word-Masking-Squad2 outputs.
-The model performs with extreme efficacy on descriptive and lengthy prompts.
-However, it can be observed to struggle with advanced mathematical prompts as it is strictly limited to segment extraction.
-Mathematical prompts can come in all sorts of wrangled manners, and the BERT model cannot handle expression extraction with logical implications.
-This is why the custom-tuned T5 proves to be so powerful, with all tasks.
+Google’s multitask-trained model falls extremely short in more logically involved tasks, while the ELECTRA model does poorly overall.
+Further training of the T5 model with more diverse and better quality data would allow it to be the most versatile and powerful model for our framework.
 6 Future Directions
@@
 6.0.2 LLM Engine Embellishments
-Exploring configurations regarding transformer models that classify, extract and parse information in the LLM engine will improve accuracy and inference speed.
+Exploring configurations regarding transformer models that classify, extract, and parse information in the LLM engine will improve accuracy and inference speed.
 The choice between delegating tasks to different models versus using a singular multi-task model for all tasks is open to be explored.
 6.0.3 AutoGen
-Expanding on the back-end embellishments, making use of Microsoft’s AutoGen (Wu et al.
-(2023)), Mistral’s Mixtral 8x7B mixture of experts model (Jiang et al.
-(2024)), or Ollama’s model hot-swapping to allow for multi-model agent setups may be the ultimate solution for delegating a task to the appropriate model rather than having a singular model handle all tasks.
-If an AutoGen setup can delegate a task to the best-suited LLM available in the engine, accuracy can be maximized across all tasks and inference times can be dynamic depending on the complexity of the task, rather than potentially sluggish times even for simpler tasks that don’t require larger models.
+Expanding on the back end embellishments, making use of Microsoft’s AutoGen [11], Mistral’s Mixtral 8x7B mixture of experts model [12], or Ollama’s model hot-swapping to allow for multi-model agent setups may be the ultimate solution for delegating a task to the appropriate model.
+AutoGen’s optimized mixture of experts system would pose many benefits over a manual swapping system, especially with the capability for agents to communicate with one another.
+If an AutoGen setup can delegate a task to the best-suited LLM amongst each other, accuracy can be maximized across all tasks and inference times can be further optimized.
 6.0.4 Concurrent Processing
 Further scaling would call for large-scale concurrent processing of incoming user requests.
-Whilst our backend server is currently multi-threaded and can handle many inputs simultaneously, further scaling to incorporate asynchronous processing with expanded capabilities, along with more robust load balancing and distributed computing systems, would bulletproof our framework’s servers.
+Whilst our back-end server is currently multi-threaded and can handle many inputs simultaneously, further scaling to incorporate asynchronous processing with expanded capabilities, along with more robust load balancing and distributed computing systems, would bulletproof our framework’s servers.
+7 Conclusion
+Our proposed LM-UI framework re-imagines the development process of contemporary user interface systems.
+We employ a system that automates the creation of textual semantic UI component annotations, which are supplementary to an agent-based prompting back end.
+This hybrid approach harnesses the power of LLMs to map generic, vague, and convoluted user textual input to their relevant components.
+We can further parse the input through logical entity extraction processes, creating a JSON object capable of dynamically adapting the UI.
+Our evaluations showcase the accuracy and precision of this approach, allowing us to take a large step closer to bringing what was once a science fiction dream, to reality.
+Furthermore, the proposed framework is just the beginning.
+The extensions and features we can implement atop this core framework increase the benefits substantially.
+From streamlining complex workflows in industries such as finance and healthcare to revolutionizing highly productive multitasking environments in modern workplaces, the implications are vast and worthy of exploration.
+We aim to continue our research efforts, to inspire and propagate a new wave of research and development in this direction.
+Our objective is to fully leverage the potential of LLMs in creating more adaptive, intuitive, and efficient software systems.